# RDGENAI ’25: First International Workshop on Retrieval-Driven Generative AI at CIKM 2025, November 14th, 2025, Coex, Seoul, Republic of Korea

### v1

@@ 3.2 Agentic RAG: Task Recognition Router, Graph RAG, Reflection
  • C1 (Color-Ingredient Visual Search): Processes questions related to color keywords ("red", "golden", "blue", etc.) or visual appearance ("elegant", "layered", "beautiful"). The Router detects based on the association between color and ingredients, and visual appeal.
@@
    Iterative Refinement: If the set quality threshold is not met, LLMs gradually expand the top-k value (a search parameter) to perform a re-search (top-k, top-(k+1), top-(k+2), …). This process repeats up to three times, improving the diversity and quality of candidate cocktails in each iteration compared to the previous results.
@@
    Final Selection: The result with the highest overall score is ultimately selected from a set of candidates. LLMs present the recommended result to the user along with the reason for the selection.
@@ A.1 Specific Graph Retrieval Algorithm
In this part, we described the topics covered in section 3.2. In particular, we have thoroughly covered the methods and processes for retrieving graph database, including examples.
C1: Color-Ingredient Visual Search Algorithm. We perform hybrid search combining graph structure and embeddings, focusing on the visual characteristics of cocktails. We extract the cocktail name, ingredients, category, and color from the user query. We then embed the user query to calculate cosine similarity with the image description embeddings, selecting the top-k cocktail nodes as seeds for graph exploration. This effectively reduces the overall graph search space. Next, it searches for ingredient nodes semantically associated with color keywords extracted from the user query (e.g., "red", "orange"). It calculates the similarity between the color keyword embedding and the ingredient name embedding to identify the most relevant ingredient, then expands the graph to include cocktails containing that ingredient. After integrating the initial seed and the expanded cocktails, we remove duplicates and recalculate the similarity between the image description embeddings of all candidates and the user query. Finally, we sort them by similarity and select the top-k cocktails. By leveraging both ingredient and image description information, we recommend cocktails that best match the user’s visual intent.
C2: Glass Type-Ingredient Matching. We explore the graph structure centered around glass type and ingredients. From user queries explicitly mentioning a glass type, extract the GlassType, Cocktail, Ingredients, and Category. First, filter Cocktail nodes matching the extracted GlassType. Then perform Ingredient matching. At this step, search for cocktails matching all ingredients. If insufficient results are found, gradually relax the ingredient condition by excluding one ingredient at a time, starting with the last ingredient extracted from the user query. This iteration continues until k candidate cocktails are secured. Finally, the final cocktail is selected by sorting the similarity scores between the extracted cocktail’s image description embedding and the user query in descending order. This approach provides results aligned with the user intent to explore cocktails with similar ingredients within the same glass type.
@@
C4: Material-based Similar Recipe Cocktail Recommendation. It comprehensively analyzes graph-based similarity and ingredient complexity to recommend alternative cocktail recipes. First, it extracts the target cocktail and ingredients from the user query as seeds. Prioritize exploring Cocktail nodes matching the seed cocktail extracted from the database. If the seed cocktail cannot be found in the user query, embed the entire query and return cocktails with high cosine similarity to the name_embedding as potential seeds. Then, explore other cocktails sharing ingredients identical to those in the seed cocktail. For each cocktail, the number of ingredients shared with the seed cocktail is calculated, while also determining the total number of ingredients in that cocktail. During the complexity analysis phase, only cocktails where the difference in ingredient count between the seed cocktail and the candidate cocktail is within m are selected. Additionally, only cocktails sharing at least x ingredients are retained. The final ranking is sorted in descending order based on the number of shared ingredients. In case of a tie, the top-k simpler cocktails with fewer ingredients are selected. This graph relationship enables the recommendation of alternative cocktails with similar ingredient compositions and appropriate complexity. C4 focuses on structural similarity rather than semantic similarity, providing practical recommendations that consider actual manufacturability and taste similarity.
@@ A.3 Detailed Description of Evaluation Metrics
For each item, LLMs instructed to evaluate on a scale of 1 to 5 points. 1 point was set as Strongly Disagree, 2 points as Disagree, 3 points as Neutral, 4 points as Agree, and 5 points as Strongly Agree. In other words, 1 point is the lowest score, and 5 points is the highest score. These evaluation items were applied identically to the Human Evaluation process.

### v2

# Proceedings of RDGENAI ’25: First International Workshop on Retrieval-Driven Generative AI at CIKM 2025, November 14th, 2025, Coex, Seoul, Republic of Korea

@@ 3.2 Agentic RAG: Task Recognition Router, Graph RAG, Reflection
  • C1 (Color-Ingredient Visual Search): Processes questions related to color keywords ("red", "blue", etc.) or visual appearance ("elegant", "layered"). The Router detects based on the association between color and ingredients, and visual appeal.
@@
    Iterative Refinement: If the set quality threshold is not met, LLMs gradually expand the top-k value (a search parameter) to perform a re-search (top-k, top-(k+1), top-(k+2), …). This process repeats up to three times, improving the diversity and quality of candidate cocktails.
@@
    Final Selection: On iteration if any top-k satisfies the threshold, the process stops immediately and that candidate set is selected as the final output. If no candidate set exceeds the threshold after all iterations, the system selects the set with the highest average score. If there are multiple sets with the same average, the one with the larger k is chosen, so that the model can generate its answer based on richer contextual information.
@@ 5 Conclusion
Declaration on Generative AI
During the preparation of this work, the author(s) used DeepL and ChatGPT for translation, grammar, and spelling check. After using this tool/service, the author(s) reviewed and edited the content as needed and take(s) full responsibility for the publication’s content.
@@ A.1 Specific Graph Retrieval Algorithm
C1: Color-Ingredient Visual Search Algorithm. C1 perform hybrid search combining graph structure and embeddings, focusing on the visual characteristics of cocktails. We extract the cocktail name, ingredients, category, and color from the user query. Next, it searches for ingredient nodes semantically associated with color keywords extracted from the user query (e.g., "red", "orange"). It calculates the similarity between the color keyword embedding and the ingredient name embedding to identify the most relevant ingredient, then locate cocktails that are jointly referenced by each attributes. For each candidate cocktail, we calculate the cosine similarity between its image description embedding and the user query embedding. Finally, all cocktails are ranked from highest to lowest similarity, and the top-k results are selected as the final recommendations. By leveraging both ingredient and image description information, we recommend cocktails that best match the user’s visual intent.
C2: Glass Type-Ingredient Matching. C2 explore related cocktails centered around the glass type and ingredients specified by the user. From user queries explicitly mentioning a glass type, extract the GlassType, Cocktail, Ingredients, and Category. First, filter Cocktail nodes matching the extracted GlassType. Then perform Ingredient matching. At this step, search for cocktails matching all ingredients. If insufficient results are found, gradually relax the ingredient condition by excluding one ingredient at a time, starting with the last ingredient extracted from the user query. This iteration continues until k candidate cocktails are secured. Finally, the final cocktail is selected by sorting the similarity scores between the extracted cocktail’s image description embedding and the user query in descending order. This approach provides results aligned with the user intent to explore cocktails with similar ingredients within the same glass type.
@@
C4: Material-based Similar Recipe Cocktail Recommendation. C4 comprehensively analyzes graph-based similarity and ingredient complexity to recommend alternative cocktail recipes. First, it extracts the target cocktail and ingredients from the user query as seeds. Prioritize exploring Cocktail nodes matching the seed cocktail extracted from the database. If the seed cocktail cannot be found in the user query, embed the entire query and return cocktails with high cosine similarity to the name_embedding as potential seeds. Then, explore other cocktails sharing ingredients identical to those in the seed cocktail. For each cocktail, the number of ingredients shared with the seed cocktail is calculated, while also determining the total number of ingredients in that cocktail. During the complexity analysis phase, only cocktails where the difference in ingredient count between the seed cocktail and the candidate cocktail is within m are selected. Additionally, only cocktails sharing at least x ingredients are retained. The final ranking is sorted in descending order based on the number of shared ingredients. In case of a tie, the top-k simpler cocktails with fewer ingredients are selected. This graph relationship enables the recommendation of alternative cocktails with similar ingredient compositions and appropriate complexity. C4 focuses on structural similarity rather than semantic similarity, providing practical recommendations that consider actual manufacturability and taste similarity.
@@ A.3 Detailed Description of Evaluation Metrics
For each item, LLMs instructed to evaluate on a scale of 1 to 5 points. 1 point was set as Strongly Disagree, 2 points as Disagree, 3 points as Neutral, 4 points as Agree, and 5 points as Strongly Agree.